Asymmetric Fermi Surfaces for Magnetic Schrödinger Operators

Horst Knorrer ETH Zürich/Institute for Advanced Study

Consider Schrödinger operators with peridic magnetic field having zero flux through a fundamental cell of the period lattice. We show that, for a generic small magnetic field and a generic Fermi energy, the corresponding Fermi surface is convex and not invariant under inversion in any point